OpenStudy (anonymous):

x^2+9xy-36y^2 factor

OpenStudy (cwrw238):

you need to find 2 terms whose product is -36 and when added give -9 can you think of them?

OpenStudy (cwrw238):

* 2 numbers

OpenStudy (anonymous):

9 and 4

OpenStudy (cwrw238):

the factors will be of the form ( x +/- _y)(x +/1 _y) where the numbers replace the blanks

OpenStudy (cwrw238):

no afraid not 9*4 = 36 ( not -36 ) and 9+4 = 13

OpenStudy (anonymous):

so what is it

OpenStudy (anonymous):

-9 and 4?

OpenStudy (cwrw238):

no - they are correct for -36 but -9 + 4 = -5 think of 2 different numbers

OpenStudy (anonymous):

(x-9) x+12)?

OpenStudy (cwrw238):

these 2 are nearly correct -12 and 3 can you see whats wrong?

OpenStudy (anonymous):

(x-3) (x+12)?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

but theres x and y involved correct is ( x + 12y(x - 3y)

OpenStudy (cwrw238):

yw